# The Lambda 1405 at the $SU(3)$ point in lattice QCD

### v1

@@ 1 Introduction
The $\Lambda$(1405) is a resonance found in the $\pi\Sigma$ channel with a mass of 1405 MeV and a with of 50.5 MeV, and it sits just below the $\overline{K}N$ threshold. It is categorized as a four star resonance in the PDG [1] with spin-parity $J^{P}=\frac{1}{2}^{-}$; strangeness S=1S=-1; and Isospin $I=1$. Thus, the expected quark content of this resonance is $(uds)$, however, it is known that the naive quark model is not fully capable of accounting for this resonance [2]. Most notably, the predicted mass for these quantum numbers is larger than what is actually observed. A potential explanation for the $\Lambda(1405)$ comes from Unitary Chiral Perturbation Theory (UChPT) [3]. In this framework the $\Lambda$(1405) is explained as a two pole resonance originating in the baryon meson interaction at the flavor $SU(3)$ symmetric point.
@@
To study the $\Lambda$(1405) with UChPT, one needs some external input, in order to fix the low energy constants. This input can come in the form of experimental data, lattice QCD energy levels or a combination of both [4, 5]. The first studies in lattice QCD into the $\Lambda$(1405) looked at isolating the lowest laying finite volume spectrum using single baryon three quark operators [6, 7, 8, 9, 10, 11, 12, 13, 14]. More recently there have been studies using a coupled channel scattering analysis with Baryon-Meson operators in gauge configurations away from the $SU(3)$ limit and at a pion mass of $M_{\pi}\approx 200$ MeV [15, 16]. The results in these papers support the UChPT picture of a two pole resonance. A separate study, using the HAL QCD method has investigated the $S$-wave Meson-Baryon interaction in the singlet and octet channels in the $SU(3)$ limit [17, 18]. For a review on the $\Lambda$(1405) considering theoretical and experimental developments see [19].
Because the origin of the two pole structure is related to the strength of the attractive forces of the decomposed interaction, looking directly at the energy levels of the different irreducible representations can help understand the nature of the $\Lambda$ (1405) and its chiral dynamics. The $SU(3)$ point can be directly accessed with lattice QCD simulations, providing a way to study the origin of the two-pole structure directly at the symmetric point. In UChPT at the leading order, the poles coming from the octets are degenerate in the $SU(3)$ limit. This accidental symmetry is broken once once incorporates next-to-leading order (NLO) terms [20]. Furthermore the trajectories that the poles follow connecting the $SU(3)$ limit with the physical point vary a lot. To study these properties, we use Baryon-Meson interpolation operators that belong to the attractive irreducible representations to measure their corresponding energy levels. We look at the location of these energy levels and the relative ordering among them. We also look at whether the energy levels coming from the octets are non-degenerate. It is also possible to use these energy levels as input for UChPT and study the trajectories of the poles the starting at the $SU(3)$ point.
@@ 3 Calculation details and lattice setup
Our measurements are performed in gauge configurations generated by the CLS collaboration [22]. These gauges are generated with a Clover-Wilson action with $N_{f}=2+1$ dynamical fermions with $m_{u}=m_{d}=m_{s}$ the details are given in Table 1 . The lattice spacing is $a\approx 0.086\text{fm}$, $V=48^{3}\times 96$ and $m_{\pi}L\approx 14.5$ which is very large and allows us to have the exponentially suppressed finite volume effects under control. In total for this work we have used 400 gauge configurations.
@@ 4 Results
Then we look at the ordering of the energy levels and we can see that the singlet is non-degenerate with the octets and that it has lower energy ($E_{1}<E_{8},E_{8^{\prime}}$). This agrees with previous lattice results [18] and with the predictions of UChPT [3]. Next, we want to look at the splitting between the different states so we perform correlated differences to account for the correlations in the uncertainties. The result is shown in Fig. 2 on the right plot. We confirm then that the singlet is non degenerate with the octets and we also find that the difference of the two octets is non zero (within one sigma), meaning that they are non degenerate, which agrees with the results from NLO UChPT [5].

### v2

@@ 1 Introduction
The $\Lambda$(1405) is a resonance found in the $\pi\Sigma$ channel with a mass of 1405 MeV and a with of 50.5 MeV, and it sits just below the $\overline{K}N$ threshold. It is categorized as a four star resonance in the PDG [1] with spin-parity $J^{P}=\frac{1}{2}^{-}$; strangeness S=1S=-1; and Isospin $I=0$. Thus, the expected quark content of this resonance is $(uds)$, however, it is known that the naive quark model is not fully capable of accounting for this resonance [2]. Most notably, the predicted mass for these quantum numbers is larger than what is actually observed. A potential explanation for the $\Lambda(1405)$ comes from Unitary Chiral Perturbation Theory (UChPT) [3]. In this framework the $\Lambda$(1405) is explained as a two pole resonance originating in the baryon meson interaction at the flavor $SU(3)$ symmetric point.
@@
To study the $\Lambda$(1405) with UChPT, one needs some external input, in order to fix the low energy constants. This input can come in the form of experimental data, lattice QCD energy levels or a combination of both [5]. The first studies in lattice QCD into the $\Lambda$(1405) looked at isolating the lowest laying finite volume spectrum using single baryon three quark operators [6, 7, 8, 9, 10, 11, 12, 13, 14]. More recently there have been studies using a coupled channel scattering analysis with Baryon-Meson operators in gauge configurations away from the $SU(3)$ limit and at a pion mass of $M_{\pi}\approx 200$ MeV [15, 16]. The results in these papers support the UChPT picture of a two pole resonance. A separate study, using the HAL QCD method has investigated the $S$-wave Meson-Baryon interaction in the singlet and octet channels in the $SU(3)$ limit [17, 18]. For a review on the $\Lambda$(1405) considering theoretical and experimental developments see [19].
Because the origin of the two pole structure is related to the strength of the attractive forces of the decomposed interaction, looking directly at the energy levels of the different irreducible representations can help understand the nature of the $\Lambda$ (1405) and its chiral dynamics. The $SU(3)$ point can be directly accessed with lattice QCD simulations, providing a way to study the origin of the two-pole structure directly at the symmetric point. In UChPT at the leading order, the poles coming from the octets are degenerate in the $SU(3)$ limit. This accidental symmetry is broken once once incorporates next-to-leading order (NLO) terms [20]. Furthermore the trajectories that the poles follow connecting the $SU(3)$ limit with the physical point vary a lot from. To study these properties, we use Baryon-Meson interpolation operators that belong to the attractive irreducible representations to measure their corresponding energy levels. We look at the location of these energy levels and the relative ordering among them. We also look at whether the energy levels coming from the octets are non-degenerate. It is also possible to use these energy levels as input for UChPT and study the trajectories of the poles the starting at the $SU(3)$ point.
@@ 3 Calculation details and lattice setup
Our measurements are performed in gauge configurations generated by the CLS collaboration [22, 23]. These gauges are generated with a Clover-Wilson action with $N_{f}=2+1$ dynamical fermions with $m_{u}=m_{d}=m_{s}$. The lattice spacing is $a\approx 0.086\text{fm}$, $V=48^{3}\times 96$ and $m_{\pi}L\approx 14.5$, see Table 1 which is very large and allows us to have the exponentially suppressed finite volume effects under control. In total for this work we have used 400 gauge configurations.
@@ 4 Results
Then we look at the ordering of the energy levels and we can see that the singlet is non-degenerate with the octets and that it has lower energy ($E_{1}<E_{8},E_{8^{\prime}}$). This agrees with previous lattice results [18] and with the predictions of UChPT [4, 3]. Next, we want to look at the splitting between the different states so we perform correlated differences to account for the correlations in the uncertainties. The result is shown in Fig. 2 on the right plot. We confirm then that the singlet is non degenerate with the octets and we also find that the difference of the two octets is non zero (within one sigma), meaning that they are non degenerate, which agrees with the results from NLO UChPT [5].